the work done when 6.5 g of zinc reacts with dil Hcl iss an open container at 298 K

Grade:12

2 Answers

Vikas TU
7 years ago
Dear Madhumitha,
I have already answered this eqn. plz check,
moles of zinc = 6.5/65 = 0.1mol
Wext. = – P*V = -1*V = -V => -nRT/P
=> -0.1*0.0821*298/1
=> -2.44658 Joules
Priyanshi jain
6 years ago
No. Of moles= 6.5/65= 0.1mol. We know that Work in a chemical reaction can be calculated by this formula , W = -ngRT. W= -0.1mol*2Cal/K/Mol*298K =59.6 Calorie.